iSuckAtMath101:

Show all work to identify the asymptotes and state the end behavior of the function f(x) = 6x/x-36 (x-36 are both under 6x)

iSuckAtMath101:

hi mr. jigglypuff314

Angle:

https://www.ck12.org/book/ck-12-algebra-i-concepts/section/12.3/ if the x values in the numerator and the x values in the denominator have the same power then the coefficients of those x values (as a fraction) is the asymptote

iSuckAtMath101:

the asymptotes are x=36 and y = 6 right?

Angle:

correct

iSuckAtMath101:

ok but how do u do that

iSuckAtMath101:

like working it out

Angle:

Rule 1: you cannot have a value of zero in the denominator when x=36 in the denominator, what happens?

iSuckAtMath101:

36 is the denominator?

iSuckAtMath101:

but its x-36 not x=36

Angle:

nono \(\frac{6x}{x-36}\) what happens when x is equal to 36?

iSuckAtMath101:

oh then a 0 would be in the denominator

iSuckAtMath101:

cs 36-36

Angle:

yes, so that is why there is an asymptote at x=36 it is because having 0 in the denominator is an illegal move in math

iSuckAtMath101:

ah

iSuckAtMath101:

ok

iSuckAtMath101:

so then what abt the 6 asymptote?

Angle:

@angle wrote:
https://www.ck12.org/book/ck-12-algebra-i-concepts/section/12.3/ if the x values in the numerator and the x values in the denominator have the same power then the coefficients of those x values (as a fraction) is the asymptote

iSuckAtMath101:

but there isnt any exponets

Angle:

FOR EXAMPLE if you have \(f(x) = \frac{3x}{2x-7}\) then the asymptote would be 3/2 because 3 is the coefficient of x in the numerator and 2 is the coefficient of x in the denominator

Angle:

x by itself has an exponent of 1

iSuckAtMath101:

OHH

iSuckAtMath101:

OHHH

iSuckAtMath101:

so what about the end behavior now?

Angle:

plug in numbers close to those values and identify what happens like, if you plug in x = 36.5 and compared it to x = 35.5

iSuckAtMath101:

6*36.5/36.5-36?

Angle:

yes

iSuckAtMath101:

6*36.5 = 219

iSuckAtMath101:

36.5 - 36 = .5

iSuckAtMath101:

219 / .5 = 438

Angle:

so x=36.5 has a positive value on the graph on the other hand x = 35.5 has a negative value on the graph

Angle:



iSuckAtMath101:

ah

iSuckAtMath101:

so i would put

iSuckAtMath101:

that the end behavior is

iSuckAtMath101:

positive?

iSuckAtMath101:

i thought it was asking like yk who they do, as x approaches "blank" f(x) increases/decreases

Angle:

you see, this is the problem -> as x approaches 36, it is decreasing on the left and increasing on the right the term here is that it "diverges" but I'm not sure if that is what your teaching is asking for

iSuckAtMath101:

oh

iSuckAtMath101:

i could say that it inreases on the right and decreases on the left?

Angle:

sure, x goes to negative infinity as you approach x=36 from the left x goes to positive infinity as you approach x=36 from the right

iSuckAtMath101:

ok

Angle:

oh wait, I ment "y"

Angle:

y goes to negative infinity as you approach x=36 from the left y goes to positive infinity as you approach x=36 from the right

iSuckAtMath101:

ah ok

iSuckAtMath101:

thanks so much

iSuckAtMath101:

u were a very big help

Angle:

no problemo glad I could help it's been a while and tbh I just did a bit of googling lol

iSuckAtMath101:

LOLL